Monday July 12
The day shift began with Operations monitoring store 3650, with the Antiproton Source (Pbar) stacking, with Operations monitoring the Recycler stash, and with MiniBooNE taking beam.
Operations setup the Recycler for stashing at 9:58 AM. This halted stacking and all Switchyard (SY) beam.
A RF expert reported at 10:11 AM that the Main Injector (MI) RF station 13 (MIRF13) was back in operation. MI had its full complement of eighteen RF stations.
A cryo system expert reported at 10:27 AM that the Tevatron (TeV) sector E1 wet engine had a failing bearing.
The Recycler finished its stash at 11:17 AM. Pbar resumed stacking.
Operations began establishing beam to the Meson MTest experiment at 1:11 PM.
At 2:14 PM, Operations halted stacking so RF experts could work on MIRF9.
At 2:16 PM, the Meson sector MS2 LCW system alarmed on high supply temperature. A Meson LCW technician investigated.
The MIRF9 work finished at 2:50 PM. Pbar resumed stacking.
A MI expert began a slipstacking study at 2:55 PM. MiniBooNE was off for much of this study.
At 3:56 PM, Operations reported that a SY SEPTA (FSEPV) had turned off for the third time today. So far it had reset itself each time.
Operations reported at 4:44 PM that MI slipstacking studies had finished.
At 4:57 PM, Operations reported that MIRF13 had tripped off numerous times in a row. Operators investigated. An expert brought the station back on line at 5:16 PM.
MiniBooNE resumed taking beam at 4:57 PM
An expert began changing the control module for FSEPV at 5:05 PM. SY beam resumed at 5:26 PM.
At 6:42 PM, FSEPV turned itself off. It turned itself back on.
MTest experimenters said that they were finished with beam at 7:45 PM.
At 8:29 PM, Operations reported that they were having problems keeping MiniBooNE up and running. Operators tuned the beamline manually, made some other adjustments, and had MiniBooNE running normally about hour later.
Tuesday July 14
The midnight shift began with Operations monitoring store 3650 and the stash, with Pbar stacking, and with MiniBooNE taking beam.
Operations established beam to the Meson MCenter experiment at 1:27 AM.
TeV experts conducted a study period from 5:14 AM to 6:09 AM.
Operations terminated store 3650 at 6:11 AM.
Cryo system techs began the TeV sector E1 wet engine bearing replacement at 6:26 AM. Operators began TeV tune tracker studies.
A tech replaced the power supply for the SY FSEPV at 7:50 AM.
The Cryo Coordinator reported at 8:53 AM that wet engine maintenance was needed at B4, D2, A1, and D3. The Run Coordinator gave permission for the work to proceed.
Operators prepared MiniBooNE for an access at 12:02 PM.
A Linac expert tuned up the Neutron Therapy Facility (NTF) at 12:35 PM for some short runs.
Cryo system techs reported that they were finished with E1 wet engine work at 1:24 PM.
An expert began a one-hour Pbar Debuncher cooling measurement study at 1:34 PM.
NTF studies finished at 3:28 PM.
Operations established beam to MiniBooNE and Pbar at 3:35 PM.
Operations put the TeV through a wet squeeze at 4:10 PM.
Operations put the TeV into shot setup at 4:34 PM and established store 3653 at 6:32 PM. Its initial luminosity was 75.2E30 Pbar resumed stacking.
Operations began stashing antiprotons at 9:40 PM.
Pbar resumed stacking at 10:48 PM.
Wednesday July 14
The midnight shift began with Operations monitoring store 3653 and the stash, with Pbar stacking, and with MiniBooNE taking beam.
At 5:53 AM, a TeV RF station (TRF8) tripped off. Operations bypassed the station.
At 7:04 AM, a Booster RF station (BRF8) tripped off and wouldn't reset. Pbar resumed stacking a few minutes later. An RF expert investigated.
Operations turned off the Booster beam while an RF expert disconnected BRF8. Beam resumed about twelve minutes later.
Plans for Wednesday Morning
The plans for today are to continue with stacking and storing.
Accelerator Update Archive
For Tevatron luminosity charts and the current status of Fermilab's
accelerators and detectors (live!), please go to Fermilab Now
Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: firstname.lastname@example.org.